what is the equation of a plane passing through (0,-2,3) and containing the x-axis??

Grade:12

2 Answers

jathin
8 years ago
by assuming the point 0i-2j+3k to be perpendicular and an arbitrary point lying on xaxis i and a point p on the plane xi +yj +zk the answer is x=1
PASUPULETI GURU MAHESH
8 years ago
p { margin-bottom: 0.25cm; line-height: 120%; }

GIVEN POINT(0,-2,3). WHERE X=0

THEN THE POINT IN THE YZ-PLANE WHERE Y=-2;Z=3

ONE LINE IS PASSINNG THROUGH THE POINT IN YZ-PLANE

 

THEN SLOPE m=3/-2 ;POINT (-2,3)

WE KNOW LINE EQATION (Z-Z1)=m(Y-Y1)

THEN (Z-3)=3/-2(Y-(-2))

3Y+2Z=0

THERE FORE LINE EQUATION IS 3Y+2Z=0